When one atom of the radioactive isotope 239Pu decomposes in a nuclear reactor, it produces a photon of gamma radiation with a wavelength of 1.590 x 10-13m.

a) How many kilojoules of energy would be released if one gram of 239Pu decomposed?

b) Use the de Broglie equation to calculate the wavelength of one atom of 239Pu travelling at 70.0% the speed of light.
Answer #1

Here,

wavelength of photon = 1.59 *10^-13 m

a)

for one gram of 239 Pu

Number of moles of Pu , n = 1/239

number of Pu atoms , N = (1/239) * 6.022 *10^23

Now ,

total energy produced = number of atoms * energy of 1 photon

total energy produced = (1/239) * 6.022 *10^23 * (6.626 *10^-34 * 3 *10^8 /(1.59 *10^-13))

total energy produced = 3.15 *10^9 J

total energy produced = 3.15 *10^6 kJ

the total energy produced is 3.15 *10^6 kJ

b)

speed , v = 0.7c

for the wavelength is

wavelength = h/(m * v)

wavelength = h *sqrt(1 - (v/c)^2)/(m * v)

wavelength = 6.626 *10^-34 * sqrt(1 - 0.7^2)/(0.7 * 3 *10^8 * 239 * 1.66 *10^-27)

wavelength = 5.68 *10^-18 m

the de broglie wavelength is 5.68 *10^-18 m